# Xiangqi Wang1, Xiuying Chen2, Mohamed Elhoseiny3, Xiangliang Zhang1,†

### v1

@@ 4.5 Human Evaluation
We conducted human evaluations in two aspects: the effectiveness of guided description generation and the alignment between questions and reference answers. See the Appendix D for details on the human evaluation. We ultimately represented the results of the human evaluation using the alignment rate (the proportion of aligned samples out of the total).
Table 4: Alignment rate of guided description generation.
Task Easy Medium Hard
Before Guide 94.32% 82.30% 77.14%
After Guide 95.20% 88.13% 84.55%
$\Delta$$\uparrow$ 0.88% 5.83% 7.41%
Guided description generation. We developed description generation guidelines for each fine-grained aspect to reduce vagueness in image descriptions, ensuring better alignment with themes and preventing discrepancies. A human evaluation showed that these guidelines significantly improved question-answer alignment, especially in more challenging tasks, as shown in LABEL:Q&A.
Question-Answer alignment. After implementing the generation guidelines, we conducted a human evaluation to assess the accuracy of the questions and answers generated by the examiner model. As shown in LABEL:Q&A, the evaluation resulted in high scores, confirming the model’s effectiveness in producing well-aligned question-answer pairs for image-based tasks.
4.6 Position Bias
Figure 9: Comparison of answer distribution under position bias conditions. Correct answers at A or D v.s. correct answers are evenly distributed at A,B,C,D.
Since the reference answers generated by LLMs tend to cluster around option A, we manually set the correct options to be evenly distributed. To investigate the necessity of this approach, we conducted experiments to examine potential position bias (Zheng et al., 2023). We evaluated scenarios where all correct answers were placed in either options A and D, comparing the resulting scores with the evenly distributed case (i.e., 25% for each option), as shown in Figure 9. The deviation rate was calculated using the following formula: $R=\frac{S_{X}-S_{U}}{S_{U}}$, where $S_{X}$ is the model score for condition $X$ (either $A$ or $D$), and $S_{U}$ is the score for the scenario when options are evenly distributed.
The position bias becomes more evident with increasing question difficulty. For instance, at the hard level, GLM-4V showed a significant bias, with deviation rates of $R_{A}=-19\%$ and $R_{D}=-8\%$, suggesting a notable bias when correct answers were concentrated in options A or D, compared to the uniform distribution scenario. Thus, our approach of manually setting an even distribution of answers to avoid position bias is justified and necessary.
@@ 5 Conclusion
In this work, we introduce AutoBench-V, a fully automated framework designed for benchmarking LVLMs. The framework integrates a series of innovative modules that ensure diversity and reliability in dataset generation, as well as impartiality in model evaluation. Through extensive experiments, we have demonstrated the robustness and unbiased nature of the evaluation process facilitated by AutoBench-V. The insights gleaned from our research provide a solid foundation for future investigations in this field.
Appendix A Details of Experiment Setting
@@
Computing Resource. For our experiments, all open-source vision-language model inferences were performed locally using NVIDIA GeForce RTX 4090 GPU with 24GB VRAM.
Alignment Evaluation. Inspired by tifa (Hu et al., 2023), we generated consistency tests for images across 12 aspects:object, human, animal, food, activity, attribute, counting, color, material, spatial, location, shape, other. For details on the score distribution of the consistency tests without threshold, please refer to Table 5.
Table 5: Alignment score ($S_{i}^{d}$) distribution without setting threshold.
[…]
Spatial Understanding.
Easy 0.51% 2.54% 7.61% 89.34%
Medium 0.53% 3.17% 19.05% 77.25%
Hard 1.03% 5.64% 16.92% 76.40%
Basic Understanding.
Easy 1.94% 2.90% 14.01% 81.16%
Medium 3.38% 5.80% 14.98% 75.85%
Hard 2.86% 8.57% 24.76% 63.81%
Semantic Understanding.
Easy 0.99% 3.94% 8.37% 86.70%
Medium 0.49% 4.39% 19.02% 76.10%
Hard 1.93% 4.35% 17.39% 76.33%
Reasoning Capacity.
Easy 1.93% 3.86% 14.49% 79.71%
Medium 2.53% 6.06% 23.23% 68.18%
Hard 6.22% 11.92% 28.50% 53.37%
Atmospheric Understanding.
Easy 0.48% 2.41% 11.11% 85.99%
Medium 1.38% 2.76% 17.97% 77.88%
Hard 1.79% 6.55% 24.40% 67.26%
Figure 10: Visualization of image topic words. Topic words are converted into vectors using bge-large-en-v1.5 (Xiao et al., 2024b), then perform dimensionality reduction via t-SNE (Van der Maaten & Hinton, 2008). Topic word distribution without semantic graph (a)(c) and with semantic graph (b)(d). It can be seen that with the semantic graph the diversity of topic words increases and the over-reliance on high-degree words is reduced.
@@
Figure 12: Topic words visulization without using semantic graph under basic undertstanding.
Table 6: Performance (Accuracy) details of all models on five user inputs and three difficulty levels.
[…]
Basic. Spatial. Seman. Reason. Atmos. Average
Easy
GPT-4o 90.18% 86.09% 93.81% 88.13% 93.94% 90.43%
GPT-4o mini 90.18% 81.28% 91.24% 81.92% 95.45% 88.01%
Gemini-1.5-Flash 89.29% 81.82% 91.19% 85.31% 92.75% 88.07%
Claude-3.5-Sonnet 91.07% 83.96% 91.75% 85.31% 94.33% 89.28%
Claude-3-Haiku 89.29% 80.21% 90.72% 82.49% 91.41% 86.82%
GLM-4v 91.96% 83.96% 87.01% 92.78% 96.45% 90.43%
Qwen2-VL 90.18% 81.82% 92.27% 87.57% 96.00% 89.57%
Medium
GPT-4o 76.87% 72.25% 83.64% 81.95% 84.35% 79.81%
GPT-4o mini 76.19% 67.26% 79.09% 77.56% 84.78% 76.98%
Gemini-1.5-Flash 73.47% 66.96% 78.18% 70.44% 84.14% 74.64%
Claude-3.5-Sonnet 75.51% 67.84% 77.73% 74.63% 81.74% 75.49%
Claude-3-Haiku 72.11% 64.76% 78.90% 67.31% 83.91% 73.40%
GLM-4v 74.83% 74.89% 79.91% 74.63% 82.17% 77.29%
Qwen2-VL 82.31% 74.01% 80.45% 73.17% 85.21% 79.03%
Hard
GPT-4o 69.12% 68.28% 79.36% 76.50% 81.82% 75.02%
GPT-4o mini 63.43% 61.23% 81.65% 69.94% 77.27% 70.70%
Gemini-1.5-Flash 66.91% 65.20% 77.10% 68.16% 76.88% 70.85%
Claude-3.5-Sonnet 59.56% 64.60% 66.97% 58.47% 68.75% 63.67%
Claude-3-Haiku 59.56% 58.15% 73.39% 71.58% 74.43% 67.42%
GLM-4v 61.03% 60.79% 66.82% 66.12% 69.89% 64.93%
Qwen2-VL 64.71% 64.76% 79.36% 71.03% 79.59% 71.89%
@@
Figure 13: We randomly sampled 75 question instances for each user input, with 25 questions from each difficulty level. We then visualized the response patterns of various models to these questions (the red means). The rows in each user input, from top to bottom, represent the response situations of Claude-3.5-sonnet, Claude-3-haiku, Gemini-1.5-flash, GLM-4v, GPT-4o, GPT-4o mini, and Qwen2-VL, respectively.
Figure 14: Visualization of all image descriptions in AutoBench-V.
@@ Appendix B Details of User Input
In this section, we provide a comprehensive overview of the levels at which we categorize user inputs based on linguistic aspects. Our goal is to offer a comprehensive and broad representation of user requirements for LVLMs. However, as it is challenging to exhaustively cover every aspect, we base our categorization on aspects derived from the literature (Li et al., 2023b; Meng et al., 2024; Liu et al., 2023b). These aspects are considered representative and comprehensive examples of the capabilities of LVLMs and other aspects like in (Chen et al., 2024b; Xu et al., 2023) can be handled in a similar manner, without requiring additional fine-tuning or adjustments, as our framework is highly extensible, allowing users to propose their own aspects as needed.
@@ B.5 Reasoning Capacity
Definition and Goal. Reasoning Capacity involves interpreting and analyzing the relationships and logical connections between different elements within an image (Zhou et al., 2024; You et al., 2023). The goal is to infer potential outcomes, understand causal relationships, and make predictions about what might happen next based on visual cues. For example, if a person is holding an umbrella and the sky is dark, reasoning capacity would suggest that it might rain soon. This level also includes understanding abstract relationships, such as social dynamics or the intent behind actions, and making judgments about what is likely or possible given the visual information.
@@ Appendix C Details of Difficulty Grading
This section describes in detail the difficulty levels for the pictures and questions used in prompts respectively. The following is the instruction guiding the examiner model to generate image descriptions and questions of varying difficulty levels.
@@ C.1 Image Description
Easy difficulty. Generate images with very simple elements, focusing on single, easily recognizable objects placed against a plain or neutral background. The descriptions should be straightforward and unambiguous, e.g., “a red apple on a white background.” The focus is on clarity and simplicity, with minimal detail or interaction.
Medium difficulty. Introduce scenes where the required elements interact with their environment naturally but uncomplicatedly. The setting may include multiple common objects and a familiar context, but the composition remains clear and not overly complex, e.g., “a cup on a table in a well-lit kitchen.” The background and context are present but not overwhelming, and there are no intricate details or unusual perspectives.
Hard difficulty. Craft descriptions that involve multiple elements interacting with each other, set in a more complex environment. Use varied perspectives, detailed textures, or lighting conditions that add layers of difficulty, e.g., “a reflection of a cat looking out of a rain-soaked window, with a cityscape in the background at dusk.” The focus is on creating a rich and intricate scene that challenges the model’s ability to render interactions, depth, and subtleties in lighting and perspective.
Moreover, we standardized the description of the observer’s perspective in the image description to prevent directional issues from causing confusion. For instance, ambiguities could arise in interpreting relative directions such as left and right, as these can vary significantly depending on the observer’s viewpoint.
@@ D.2 Human Evaluation Guidelines
Appendix E Error Study
Through extensive experimental analysis, we have categorized the common problems encountered by LVLMs in VQA tasks into two main types: image comprehension errors and image reasoning errors. Regarding the first category, LVLMs often fail to truly understand the details in an image. For instance, in Figure 15, the model failed to notice the kite flying in the upper right corner of the image and incorrectly identified a non-existent action of a squirrel climbing a tree. In Figure 17, the model mistakenly perceived the firefighter as holding the crowbar with only one hand instead of two. In Figure 18, the model failed to recognize that the apple’s stem was slightly tilted to the left. These errors demonstrate the model’s inadequate comprehension of image details.
Image reasoning errors occur when models accurately perceive the image content but falter in their reasoning process, leading to incorrect answers. For instance, in Figure 16, the model correctly recognizes that the image does not depict a cheering crowd. However, during subsequent reasoning, it convinces itself otherwise, ultimately selecting option C while neglecting to analyze other choices in its explanation. Figure 19 exemplifies a similar issue: the model correctly identifies that the child is wearing a yellow shirt but after mentioning the red of the kite it erroneously selects B. Red is its final answer. These examples highlight a disconnect between visual perception and logical reasoning in LVLMs, where initial accurate observations can be overridden by flawed deductive processes.
Figure 15: Error study of Claude-3-Haiku under basic understanding.
Figure 16: Error study of Claude-3.5-Sonnet under spatial understanding.
Figure 17: Error study of Qwen2-VL under semantic understanding.
Figure 18: Error study of GLM-4v under reasoning capacity.
Figure 19: Error study of Gemini-1.5-Flash under atmospheric understanding.

### v3

@@ 4.4 Human Evaluation
We conducted human evaluations to assess the alignment between the generated questions and the reference answers. The evaluations were performed using a metric we refer to as the alignment rate, defined as the proportion of aligned samples out of the total. Detailed procedures for the human evaluation can be found in Appendix D.
The results show that the model achieves high alignment across difficulty levels: 95.20% for Easy tasks, 88.13% for Medium, and 84.55% for Hard. This demonstrates the model’s ability to generate well-aligned question-answer pairs, excelling in simpler tasks while maintaining strong performance under higher complexity, highlighting its effectiveness in producing accurate, contextually appropriate outputs for visual tasks.
@@ 5 Conclusion
In this work, we introduce AutoBench-V, an automated framework designed to benchmark LVLM. It integrates a series of innovative modules that ensure diversity and reliability in dataset generation and evaluation. Extensive experiments demonstrate its robustness and unbiased process, offering a solid foundation for future research.
Appendix A Details of Experiment Setting & Result
@@
Alignment Evaluation. Inspired by TIFA (Hu et al., 2023), we generated consistency tests for images across 12 aspects:object, human, animal, food, activity, attribute, counting, color, material, spatial, location, shape, other.
Figure 8: Visualization of image topic words. Topic words are converted into vectors using bge-large-en-v1.5 (Xiao et al., 2024), then perform dimensionality reduction via t-SNE (Van der Maaten and Hinton, 2008). Topic word distribution without semantic graph (a)(c) and with semantic graph (b)(d). It can be seen that with the semantic graph the diversity of topic words increases and the over-reliance on high-degree words is reduced.
@@
Figure 11: Topic words visualization without using semantic graph under basic understanding.
@@
[…]
Figure 12: Visualization of image descriptions in AutoBench-V.
@@ Appendix B Details of User Input
In this section, we provide a comprehensive overview of the levels at which we categorize user inputs based on linguistic aspects. Our goal is to offer a comprehensive and broad representation of user requirements for LVLMs. However, as it is challenging to exhaustively cover every aspect, we base our categorization on aspects derived from the literature (Li et al., 2023b; Meng et al., 2024; Liu et al., 2025). These aspects are considered representative and comprehensive examples of the capabilities of LVLMs and other aspects like in (Chen et al., 2024b; Xu et al., 2024) can be handled in a similar manner, without requiring additional fine-tuning or adjustments, as our framework is highly extensible, allowing users to propose their own aspects as needed.
@@ B.5 Reasoning Capacity
Definition and Goal. Reasoning Capacity involves interpreting and analyzing the relationships and logical connections between different elements within an image (Zhou et al., 2024a; You et al., 2023). The goal is to infer potential outcomes, understand causal relationships, and make predictions about what might happen next based on visual cues. For example, if a person is holding an umbrella and the sky is dark, reasoning capacity would suggest that it might rain soon. This level also includes understanding abstract relationships, such as social dynamics or the intent behind actions, and making judgments about what is likely or possible given the visual information.
@@ Appendix C Details of Difficulty Grading
This section describes in detail the difficulty levels for the pictures (subsection C.1) and questions (subsection C.2) used in prompts respectively. The following is the instruction guiding the examiner model to generate image descriptions and questions of varying difficulty levels. For details, see Appendix H. Moreover, the difficulty levels are defined manually and are not fixed, allowing for flexible adjustments based on the user’s specific requirements.
@@ C.1 Image Description
Easy difficulty. Focus on a single subject but with minor additional complexity to test precision in details or context. Establish baseline capability with subtle challenges, such as fine texture, simple interactions, or slight variations in lighting. A single object or entity with some basic contextual details or additional features, placed against a minimally distracting background. Test finer details (e.g., texture, reflection) while still keeping the composition simple and clear. For examples: "A perfectly polished red apple with a small leaf attached, placed on a slightly reflective white surface."
Medium difficulty. Scenes involve multiple elements or interactive settings that require nuanced spatial arrangement and accurate relationships between objects. Evaluate the ability to generate cohesive, moderately dynamic scenes with layered realism and a stronger sense of depth. Description Requirements: Include multiple objects interacting naturally in a believable environment, with more intricate details and subtle light or shadow effects. Incorporate realistic environmental elements and ensure spatial coherence, emphasizing interactions and secondary details (e.g., shadows, water splashes). For examples: "A golden retriever running on a sandy beach, splashing water as it chases a bright orange ball, with distant waves and a partly cloudy sky."
Hard difficulty. Scenes incorporate high complexity with multiple interdependent elements, challenging perspectives, or dynamic and intricate lighting or textural effects. Push the limits of rendering capability to handle advanced relationships, environmental effects, and challenging compositions. The scene must include 3-4 main elements or a combination of dynamic features, such as motion, light interplay, or atmospheric conditions, while maintaining clarity and logic. Highlight challenges such as complex reflections, dynamic light, or multi-element interactions, ensuring visual harmony and detailed textures. For examples: "A raindrop-streaked window reflecting the interior of a cozy room, with a black cat sitting on the windowsill and a glowing city skyline visible through the glass, all under the warm hues of a sunset."
@@ D.2 Human Evaluation Guidelines
Appendix E Notations
Table 6: Notation used in the methodology description.
Notation Description
$q$ User input specifying the evaluation target or focus.
$n$ Number of general aspects.
$\{A^{(g)}_{1},A^{(g)}_{2},\dots,A^{(g)}_{n}\}$ Set of $n$ general aspects, each representing a high-level capability dimension.
$A^{(g)}_{i}$ The $i$-th general aspect.
$m$ Number of fine-grained aspects under each general aspect.
$\{A^{(f)}_{i1},A^{(f)}_{i2},\dots,A^{(f)}_{im}\}$ Set of $m$ fine-grained aspects under the $i$-th general aspect.
$A^{(f)}_{ij}$ The $j$-th fine-grained aspect under the $i$-th general aspect.
$\mathcal{M}_{v}$ The LVLM used for aspect generation, description generation, and self-validation.
$\mathcal{A}$ The complete set of generated aspects, i.e., $\bigcup_{i=1}^{n}\{A^{(g)}_{i}\}\cup\{A^{(f)}_{ij}\}$.
$|\mathcal{A}|$ The cardinality of $\mathcal{A}$, with $|\mathcal{A}|=m\times n$.
$d\in\{\text{easy},\text{medium},\text{hard}\}$ Difficulty level for image description generation.
$\omega$ Number of image descriptions generated per aspect at each difficulty level.
$\{\mathcal{T}_{ij1}^{d},\dots,\mathcal{T}_{ij\omega}^{d}\}$ Set of $\omega$ image descriptions for the fine-grained aspect $A^{(f)}_{ij}$ at difficulty $d$.
$\mathcal{T}_{ij}^{d}$ A specific image description for aspect $A^{(f)}_{ij}$ at difficulty $d$.
$t_{e}$ Topic word selected at iteration $e$.
$K_{e}=\{k_{e1},k_{e2},\dots,k_{ec}\}$ Set of $|c|$ keywords related to the topic word $t_{e}$ at iteration $e$.
$G_{e}=(V_{e},E_{e})$ Semantic graph at iteration $e$, whose nodes are topic/keyword sets and edges are semantic relationships.
$S_{e}$ Combination of $V_{e-1}$, $t_{e}$, and $K_{e}$ before applying the exclusion mechanism.
$V_{e}$ Refined node set (topic words/keywords) after excluding high-degree nodes in iteration $e$.
$f(S_{e})$ A function that determines how many top-degree nodes to exclude from $S_{e}$.
$\{\mathcal{I}_{ij}^{d}\}$ Generated images corresponding to the image description(s) $\mathcal{T}_{ij}^{d}$.
$\Phi_{ij}^{d}=\{\phi_{ij1}^{d},\dots,\phi_{ijp}^{d}\}$ Set of $p$ simple VQA-style questions for self-validation of $\mathcal{I}_{ij}^{d}$.
$p$ Number of questions for each image’s self-validation.
$\mathcal{F}(\mathcal{I}_{ij}^{d},\mathcal{T}_{ij}^{d},\Phi_{ij}^{d})$ Self-validation function returning the alignment score $S_{ij}^{d}$.
$S_{ij}^{d}$ Alignment score, i.e., the proportion of correctly answered questions among $\Phi_{ij}^{d}$.
$\zeta$ Threshold for deciding whether to regenerate an image or keep it with recorded errors.
$\mathcal{E}_{ij}^{d}$ Noted error(s) if the image partially misaligns with the description ($S_{ij}^{d}<1$).